### v1

# Suppression of Gravitational-Wave Echoes in Diffeomorphism-Invariant Nonlocal Gravity

Abstract
Searches for late–time gravitational–wave echoes following compact binary mergers are often interpreted as probes of exotic near–horizon or horizonless physics. We show that a well–motivated class of ultraviolet–finite quantum gravity and modified gravity theories, those based on diffeomorphism–invariant, analytic, entire–function nonlocality, generically suppress observable echo signals. The suppression follows from Paley–Wiener bounds associated with the analyticity of the nonlocal regulator, which enforce exponential damping of high–frequency reflection amplitudes and smear sharp effective potentials in the black–hole interior. We demonstrate that for regular black holes with horizons, the standard ingoing boundary condition eliminates the cavity required for echoes, while for regular horizonless compact objects the nonlocal kernel strongly attenuates both inner and photon–sphere reflections. As a result, repeated reflections during the ringdown phase are exponentially suppressed in frequency space and washed out in the time domain. Our results imply that the absence of echoes in current gravitational–wave data is consistent with covariant nonlocal gravity theories and does not, by itself, favor classical horizons over regular or horizonless ultraviolet completions. The analyticity and Paley–Wiener bounds that enforce echo suppression are required to ensure unitarity, ghost freedom, and ultraviolet finiteness of the underlying quantum theory. As a result, the classical ringdown phenomenology already encodes quantum–gravity consistency conditions. We emphasize that the absence of echoes in such theories does not rely on quantum decoherence or stochastic effects, but instead reflects the analytic structure imposed by quantum gravity on the classical limit.
1 Introduction
The detection of gravitational waves from compact binary mergers has opened a new observational window onto the strong–field regime of gravity. In particular, the post–merger ringdown phase, dominated by quasinormal modes of the remnant compact object, provides a sensitive probe of the near–horizon structure of spacetime. Motivated by this possibility, a number of works have proposed that departures from classical black holes—such as horizonless compact objects, quantum–modified horizons, or reflective surfaces replacing the event horizon—could generate late–time, repeating echoes in the gravitational–wave signal.
Echoes arise when gravitational perturbations are repeatedly trapped between an effective potential barrier near the photon sphere and a partially reflective inner boundary. In classical general relativity, an event horizon enforces purely ingoing boundary conditions, preventing such a cavity from forming. By contrast, many phenomenological echo models assume ad hoc reflective conditions at microscopic distances outside the would–be horizon or at the surface of a compact object. The observational status of echoes remains unsettled, with current data providing no statistically compelling evidence for their existence.
Ultraviolet–complete or ultraviolet–finite theories of quantum gravity can be constructed using nonlocal operators built from analytic entire functions of the covariant d’Alembertian [1, 2, 3, 4, 5, 6, 7, 8]. These theories preserve diffeomorphism invariance, avoid additional ghostlike degrees of freedom, and render quantum loop corrections finite. The nonlocality is controlled by a fundamental scale and is encoded through form factors whose analyticity implies Paley–Wiener bounds on their Fourier transforms [11, 12].
We show that these same analytic and diffeomorphism–invariant properties have direct and robust consequences for gravitational–wave ringdown phenomenology [13, 14, 15, 16, 17]. In particular, the Paley–Wiener bounds enforce an exponential suppression of high–frequency reflection amplitudes and smear any sharp features in the effective radial potential governing perturbations. As a result, the two essential ingredients required for observable echoes—strong inner reflection and a sharply defined resonant cavity—are generically absent.
We analyze both regular black holes possessing true event horizons and regular horizonless compact objects arising in nonlocal gravity. In the former case, the presence of a horizon enforces ingoing boundary conditions that eliminate inner reflection altogether. In the latter case, regularity at the center replaces the notion of a hard reflecting surface, while the nonlocal kernel suppresses scattering from interior structure and from the photon–sphere barrier itself. In both situations, any would–be echo train is exponentially damped with increasing frequency and echo number, and is further smeared in the time domain by the intrinsic nonlocal scale.
Our results demonstrate that the absence of gravitational–wave echoes is a prediction of analytic entire–function nonlocal gravity theories. Consequently, non–detection of echoes should not be interpreted as evidence against ultraviolet completions featuring regular black holes or horizonless compact objects, but rather as a consistency check of covariant nonlocal dynamics. Ringdown observations instead constrain the nonlocal scale and the analytic structure of the gravitational form factor, providing a complementary probe of quantum gravity in the strong–field regime.
The suppression of gravitational–wave echoes derived in this work operates entirely at the level of classical wave propagation on a fixed background geometry. The modified Regge–Wheeler–type equations [18, 19, 20], the smearing of effective potentials, and the exponential damping of reflection coefficients follow from classical integro–differential equations of motion. However, the specific nonlocal structure responsible for these effects, namely, analytic entire functions of the covariant d’Alembertian obeying Paley–Wiener bounds—is not motivated by classical considerations alone. Such analyticity conditions are imposed to ensure unitarity, ghost freedom, and ultraviolet finiteness of the quantum gravitational theory. In this sense, quantum gravity is already implicitly encoded in the classical dynamics. The same analytic properties required for a consistent quantum theory manifest classically as a frequency–space filter that suppresses sharp reflections and eliminates the resonant cavity necessary for observable echoes. The resulting ringdown phenomenology reflects quantum–gravity consistency conditions, even in the absence of explicit quantum fluctuations or loop effects.
2 Regular Static Spherically Symmetric Solution
A covariant entire-function form factor $\mathcal{F}(\Box/\Lambda_{G}^{2})$ with $\mathcal{F}$ entire and $\mathcal{F}(0)=1$ can be implemented so that matter sources are nonlocally dressed, while maintaining diffeomorphism invariance [3, 4, 5]. The distributional stress–energy is replaced by an effective, smooth profile:
$T_{\mu\nu}(x)\;\longrightarrow\;T^{\rm eff}_{\mu\nu}(x):=\mathcal{F}\!\left(\frac{\Box}{\Lambda_{G}^{2}}\right)T_{\mu\nu}(x),$ (1)
where $\Box=g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}$ is the covariant d’Alembertian and $\Lambda_{G}$ is the nonlocality scale. For an exponential entire function:
$\mathcal{F}\!\left(\frac{\Box}{\Lambda_{G}^{2}}\right)=e^{\Box/\Lambda_{G}^{2}},$ (2)
the action of $\mathcal{F}$ on static sources reduces in a weakly curved, quasi-static regime to a Gaussian heat-kernel smearing in space. This is the sense in which the regulator acts as a smearing map on matter distributions.
In what follows, we derive the resulting regular black-hole and compact-object geometries in the static, spherically symmetric sector, and show explicitly the emergence of a de Sitter core and the horizon phase diagram controlled by a single dimensionless parameter.
We adopt the standard static, spherically symmetric line element:
$ds^{2}=-f(r)\,dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2},\qquad d\Omega^{2}=d\theta^{2}+\sin^{2}\theta\,d\phi^{2},$ (3)
and model the smeared source as an anisotropic but static effective fluid:
$T^{\mu}{}_{\nu}=\mathrm{diag}\!\big(-\rho(r),\,p_{r}(r),\,p_{t}(r),\,p_{t}(r)\big).$ (4)
Einstein’s equations,
$G_{\mu\nu}=8\pi G\,T_{\mu\nu},$ (5)
imply that $f(r)$ can be written in terms of the Misner–Sharp mass function $m(r)$ [21]:
$f(r)=1-\frac{2G\,m(r)}{r},\qquad m^{\prime}(r)=4\pi r^{2}\rho(r),$ (6)
where prime denotes $d/dr$. Regularity at the origin requires $m(r)\sim\mathcal{O}(r^{3})$ as $r\to 0$.
Consider a point mass $M$ localized at the origin in the distributional limit. Under the smearing map (1) with the exponential choice (2), the effective energy density is well-approximated by a Gaussian profile of width $\ell\sim\Lambda_{G}^{-1}$:
$\rho(r)=\frac{M}{(4\pi\ell^{2})^{3/2}}\exp\!\left(-\frac{r^{2}}{4\ell^{2}}\right),$ (7)
normalized so that the total mass is $M$:
$4\pi\int_{0}^{\infty}dr\,r^{2}\rho(r)=M.$ (8)
The relation between $\ell$ and $\Lambda_{G}$ depends on convention $\ell=1/\Lambda_{G}$ or $\ell=1/(\sqrt{2}\Lambda_{G}$; for the horizon phase structure only the ratio $GM/\ell$ matters.
Integrating (6) with (7) gives
$m(r)$ $=4\pi\int_{0}^{r}ds\,s^{2}\rho(s)=\frac{M}{(4\pi\ell^{2})^{3/2}}\,4\pi\int_{0}^{r}ds\,s^{2}\exp\!\left(-\frac{s^{2}}{4\ell^{2}}\right)$
$=M\left[\operatorname{erf}\!\left(\frac{r}{2\ell}\right)-\frac{r}{\sqrt{\pi}\,\ell}\exp\!\left(-\frac{r^{2}}{4\ell^{2}}\right)\right],$ (9)
where $\operatorname{erf}$ is the error function. The corresponding lapse function is given by
$f(r)=1-\frac{2GM}{r}\left[\operatorname{erf}\!\left(\frac{r}{2\ell}\right)-\frac{r}{\sqrt{\pi}\,\ell}\exp\!\left(-\frac{r^{2}}{4\ell^{2}}\right)\right].$ (10)
By construction $m(r)\to M$ as $r\to\infty$, so $f(r)\to 1-2GM/r$ and the geometry is asymptotically Schwarzschild.
Near the origin, expand (9) for $r\ll\ell$:
$m(r)=\frac{4\pi}{3}\rho_{0}\,r^{3}+\mathcal{O}(r^{5}),\qquad\rho_{0}:=\rho(0)=\frac{M}{(4\pi\ell^{2})^{3/2}}.$ (11)
Then (6) yields
$f(r)=1-\frac{2G}{r}\left(\frac{4\pi}{3}\rho_{0}r^{3}\right)+\mathcal{O}(r^{4})=1-\frac{8\pi G\rho_{0}}{3}\,r^{2}+\mathcal{O}(r^{4}).$ (12)
The core is de Sitter with effective cosmological constant:
$\Lambda_{\rm eff}=8\pi G\rho_{0},\qquad f(r)=1-\frac{\Lambda_{\rm eff}}{3}r^{2}+\mathcal{O}(r^{4}),$ (13)
and all curvature invariants remain finite at $r=0$, because $\rho_{0}<\infty$ and $f(r)$ is analytic there. In particular, the would-be Schwarzschild singularity is replaced by a smooth de Sitter core. Since the Schwarzschild singularity is spacelike, its regularization necessarily replaces it with a finite spacelike de Sitter core, preserving the causal structure of the interior while eliminating curvature divergences; the resulting core is locally static but globally lacks a preferred time direction.
Horizons occur at the positive roots of $f(r)=0$. It is convenient to introduce dimensionless variables
$x:=\frac{r}{\ell},\qquad\mu:=\frac{GM}{\ell},,$ (14)
so that (10) becomes
$f(x)=1-\frac{2\mu}{x}\left[\operatorname{erf}\!\left(\frac{x}{2}\right)-\frac{x}{\sqrt{\pi}}\,e^{-x^{2}/4}\right].$ (15)
The entire horizon structure is controlled by the single dimensionless parameter $\mu=GM/\ell$, equivalently, $M$ measured in units set by the nonlocal length $\ell$.
The transition between two horizons and no horizon occurs at an extremal configuration where $f$ has a double root. Denoting the extremal radius by $x_{\star}>0$, the conditions are given by
$f(x_{\star})=0,\qquad f^{\prime}(x_{\star})=0.$ (16)
From (15) these two equations implicitly determine the critical pair $(x_{\star},\mu_{\star})$:
$\mu_{\star}=\frac{x_{\star}}{2\left[\operatorname{erf}\!\left(\frac{x_{\star}}{2}\right)-\frac{x_{\star}}{\sqrt{\pi}}\,e^{-x_{\star}^{2}/4}\right]},\qquad 0=\left.\frac{d}{dx}\left\{\frac{1}{x}\left[\operatorname{erf}\!\left(\frac{x}{2}\right)-\frac{x}{\sqrt{\pi}}e^{-x^{2}/4}\right]\right\}\right|_{x=x_{\star}}.$ (17)
This provides a sharp, regulator-controlled critical point without introducing any additional dimensionless tuning.
As $\mu$ varies, one finds the standard three-regime pattern characteristic of these smeared-source solutions. $\mu>\mu_{\star}$: two distinct horizons $x_{-}$ and $x_{+}$ inner and outer. $\mu=\mu_{\star}$: one degenerate extremal horizon at $x_{\star}$. $\mu<\mu_{\star}$: no horizon a regular, horizonless compact object. Because $f(x)$ is smooth and the core is de Sitter, the $\mu<\mu_{\star}$ branch is nonsingular and geodesically extendible through $r=0$.
The effective stress tensor (4) must satisfy $\nabla_{\mu}T^{\mu}{}_{\nu}=0$. For the metric (3) this implies:
$p_{r}^{\prime}(r)+\frac{f^{\prime}(r)}{2f(r)}\big(\rho(r)+p_{r}(r)\big)+\frac{2}{r}\big(p_{r}(r)-p_{t}(r)\big)=0.$ (18)
A minimal consistent for smeared static sources is $T^{t}{}_{t}=T^{r}{}_{r}$, $p_{r}=-\rho$, which ensures regularity and simplifies (18) to a relation determining $p_{t}(r)$. The detailed microphysical interpretation of $p_{r},p_{t}$ depends on the underlying nonlocal completion, but the geometric conclusions above follow already from the single input, a smooth, conserved $T^{\rm eff}_{\mu\nu}$ generated by the entire-function smearing map.
The covariant entire-function regulator replaces distributional matter sources by smooth effective profiles, and for a smeared point mass leads to the closed-form metric (10). The geometry is asymptotically Schwarzschild, develops a de Sitter core (13), and exhibits a sharp horizon phase diagram controlled by the single parameter $\mu=GM/\ell$, with the extremal point determined implicitly by (16)–(17). This is the sense in which the regulator simultaneously enforces UV finiteness via nonlocal damping and geometrically resolves the classical Schwarzschild singularity in the static, spherically symmetric sector.
3 Suppression of Ringdown Echoes in Covariant Nonlocal Gravity
We consider the class of covariant nonlocal theories in which the gravitational action is built from curvature invariants and analytic entire functions of the covariant d’Alembertian $\Box\equiv g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}$. A representative form is given by
$S\;=\;\frac{1}{16\pi G}\int d^{4}x\,\sqrt{-g}\;\Big[R\;+\;R\,\mathcal{F}_{1}(\Box)\,R\;+\;R_{\mu\nu}\,\mathcal{F}_{2}(\Box)\,R^{\mu\nu}\;+\;C_{\mu\nu\rho\sigma}\,\mathcal{F}_{3}(\Box)\,C^{\mu\nu\rho\sigma}\Big],$ (19)
where $R$, $R_{\mu\nu}$ and $C_{\mu\nu\rho\sigma}$ denote the Ricci scalar, Ricci and Weyl curvature tensors, respectively. Each $\mathcal{F}_{i}(z)$ is analytic entire and chosen so that no additional propagating ghost poles are introduced. In ultraviolet–finite constructions one typically encounters an entire form factor or regulator of exponential type:
$\mathcal{F}(z)\;=\;\exp\!\Big[-\big(z/\Lambda_{G}^{2}\big)^{n}\Big],\qquad n\in\mathbb{Z}_{\geq 1},$ (20)
with nonlocality scale $\Lambda_{G}$. [^1] In Euclidean momentum variables $p_{E}$, one has $\mathcal{F}(p_{E}^{2})=\exp[-(p_{E}^{2}/\Lambda_{G}^{2})^{n}]$, giving strong ultraviolet damping.
We study linearized gravitational perturbations around a static, spherically symmetric regular remnant geometry describing the post–merger object:
$ds^{2}=-f(r)\,dt^{2}+\frac{dr^{2}}{g(r)}+r^{2}d\Omega^{2},\qquad f(r),g(r)\ \text{regular for}\ r\geq 0,$ (21)
allowing either a regular black hole with a true horizon at $r=r_{h}$, where $f(r_{h})=g(r_{h})=0$ or a regular horizonless compact object, where $f,g>0$ for all $r\geq 0$. We emphasize that our conclusions do not depend on a particular regular metric model; only smoothness and the analytic nonlocal structure are used.
Upon decomposing perturbations into parity sectors and spherical harmonics, the dynamics can be expressed in terms of a master variable $\Psi(t,r)$ of Regge–Wheeler-Zerilli type [18, 19, 20]. In local GR, Fourier modes $\Psi(t,r)=e^{-i\omega t}\Psi_{\omega}(r)$ satisfy:
$\Big[\partial_{r_{*}}^{2}+\omega^{2}-V_{\rm loc}(r)\Big]\Psi_{\omega}(r)=0,$ (22)
where $r_{*}$ is the tortoise coordinate defined by $dr_{*}/dr=1/\sqrt{f(r)g(r)}$, and $V_{\rm loc}(r)$ is the effective potential with a barrier near the photon sphere with $r_{ph}=3GM$.
In analytic nonlocal gravity, the operator $\Box$ acts nontrivially on perturbations and generates nonlocal modifications that can be represented as a radial convolution by a kernel whose Fourier transform is the form factor. A general and convenient parametrization of the frequency–domain equation is given by
$\Big(\partial_{r_{*}}^{2}+\omega^{2}\Big)\Psi_{\omega}(r_{*})\;=\;\int_{-\infty}^{+\infty}dr_{*}^{\prime}\;K_{\Lambda_{G}}(r_{*}-r_{*}^{\prime})\,V_{\rm loc}(r_{*}^{\prime})\,\Psi_{\omega}(r_{*}^{\prime})\,$, (23)
where $K_{\Lambda_{G}}$ is a smooth kernel of width $\Delta r_{*}\sim\Lambda_{G}^{-1}$ in the local proper measure, determined by $\mathcal{F}$. Equation (23) states that the nonlocality smears the potential and modifies scattering.
Defining the smeared effective potential:
$V_{\rm eff}(r_{*})\;\equiv\;\int_{-\infty}^{+\infty}dr_{*}^{\prime}\;K_{\Lambda_{G}}(r_{*}-r_{*}^{\prime})\,V_{\rm loc}(r_{*}^{\prime})\,$, (24)
one may formally rewrite (23) as
$\Big[\partial_{r_{*}}^{2}+\omega^{2}\Big]\Psi_{\omega}(r_{*})\;=\;V_{\rm eff}(r_{*})\,\Psi_{\omega}(r_{*})\;+\;\Delta_{\rm NL}[\Psi_{\omega}],$ (25)
where $\Delta_{\rm NL}$ collects non-multiplicative nonlocal terms; in the Born scattering regime the convolution form (23) is sufficient to establish robust suppression bounds.
Let the Fourier transform of the kernel be given by
$\widetilde{K}_{\Lambda_{G}(k)}\;\equiv\;\int_{-\infty}^{+\infty}dx\;e^{-ikx}K_{\Lambda_{G}}(x).$ (26)
For the Gaussian regulator $n=1$ in (20):
$\widetilde{K}_{\Lambda_{G}(k)}\;=\;\exp\!\Big[-k^{2}/\Lambda_{G}^{2}\Big],$ (27)
and for general $n$ one obtains a function that decays super–polynomially, indeed faster than any power at large $|k|$. More generally, the analyticity and exponential–type bounds on $\mathcal{F}$ imply Paley–Wiener bounds on $\widetilde{K}_{\Lambda_{G}(k)}$ [11, 12], yielding strong damping in the frequency domain and smooth kernels in the radial domain. Physically, the nonlocality acts as a band–limiting filter for scattering.
To make this explicit, consider scattering at frequency $\omega$ from a localized potential barrier. In the Born approximation, the complex reflection amplitude in the local problem (22) is given by
$R_{\rm loc}(\omega)\;\simeq\;\frac{1}{2i\omega}\int_{-\infty}^{+\infty}dr_{*}\;e^{2i\omega r_{*}}\,V_{\rm loc}(r_{*}).$ (28)
Replacing $V_{\rm loc}$ by the smeared $V_{\rm eff}=K_{\Lambda_{G}}*V_{\rm loc}$, the nonlocal reflection amplitude becomes:
$R_{\rm NL}(\omega)$ $\simeq\frac{1}{2i\omega}\int dr_{*}\;e^{2i\omega r_{*}}\,V_{\rm eff}(r_{*})$
$=\frac{1}{2i\omega}\int dr_{*}^{\prime}\;V_{\rm loc}(r_{*}^{\prime})e^{2i\omega r_{*}^{\prime}}\int dx\;K_{\Lambda_{G}(x)}e^{2i\omega x}$
$=\widetilde{K}_{\Lambda_{G}}(2\omega)\,R_{\rm loc}(\omega).$ (29)
The analytic nonlocality multiplicatively suppresses reflection through the factor $\widetilde{K}_{\Lambda_{G}}(2\omega)$. For the Gaussian example (27):
$|R_{\rm NL}(\omega)|\;\simeq\;e^{-4\omega^{2}/\Lambda_{G}^{2}}\,|R_{\rm loc}(\omega)|.$ (30)
This is the basic Paley–Wiener suppression mechanism [11, 12]. The same analyticity that yields ultraviolet finiteness forces reflection coefficients to be exponentially damped at frequencies above the nonlocal scale.
The photon–sphere barrier reflectivity $R_{\rm ph}(\omega)$ entering ringdown scattering is therefore filtered as:
$R_{\rm ph,NL}(\omega)\;\approx\;\widetilde{K}_{\Lambda_{G}}(2\omega)\,R_{\rm ph,loc}(\omega),\qquad T_{\rm ph,NL}(\omega)\;\approx\;T_{\rm ph,loc}(\omega)\ \ \text{(to leading order)}.$ (31)
A similar filtering applies to any effective interior barrier that would otherwise reflect waves back toward the photon sphere.
Echo trains are commonly modeled as repeated partial reflections between the photon–sphere barrier and an inner reflector. In frequency space one may write the outgoing waveform as:
$\Psi_{\rm out}(\omega)\;\propto\;T_{\rm ph}(\omega)\,\Psi_{\rm in}(\omega)\,\sum_{n=0}^{\infty}\Big[R_{\rm ph}(\omega)\,R_{\rm in}(\omega)\,e^{2i\omega\Delta t}\Big]^{n},$ (32)
where $R_{\rm in}(\omega)$ is the inner reflection amplitude and $\Delta t$ is the round–trip time between the photon sphere and the inner scattering region expressed in tortoise coordinate. The transfer function is given by
$\mathcal{G}(\omega)\;\equiv\;\frac{1}{1-R_{\rm ph}(\omega)\,R_{\rm in}(\omega)\,e^{2i\omega\Delta t}}.$ (33)
The existence of observable echoes requires two conditions, a non-negligible inner reflectivity $|R_{\rm in}(\omega)|\not\ll 1$ over the relevant band, and sufficiently sharp scattering phase coherence to sustain repeated returns, so that $|R_{\rm ph}R_{\rm in}|$ is not rapidly suppressed with $\omega$. We now show that analytic entire–function nonlocality generically violates both conditions.
If the remnant has a true horizon at $r=r_{h}$, then $r_{*}\to$-$\infty$ as $r\to r_{h}^{+}$ and the physical boundary condition for perturbations is purely ingoing at the horizon:
$\Psi_{\omega}(r_{*})\;\sim\;e^{-i\omega r_{*}},\qquad r_{*}\to-\infty.$ (34)
This implies the absence of an outgoing component from the interior:
$R_{\rm in}(\omega)\;\simeq\;0\qquad\Rightarrow\qquad\mathcal{G}(\omega)\approx 1.$ (35)
A regular black hole with a horizon does not support an echo cavity, irrespective of the ultraviolet completion, provided the completion does not introduce a reflective surface outside the horizon. Analytic entire–function nonlocality, being smooth and diffeomorphism invariant, does not generate such a hard surface; rather it smears curvature and sources.
For a regular horizonless compact object, $r_{*}$ remains finite throughout the interior and the appropriate inner condition is regularity at the center. For typical parity sectors one has near $r=0$:
$\Psi_{\omega}(r)\;\propto\;r^{\ell+1}\quad(r\to 0),$ (36)
which is not equivalent to imposing a reflective wall. Any effective inner reflection arises instead from smooth interior structure turning points in $V_{\rm eff}$, and is therefore intrinsically sensitive to the nonlocal smearing.
Applying the same reasoning as in (29), the inner reflectivity is filtered:
$R_{\rm in,NL}(\omega)\;\approx\;\widetilde{K}_{\Lambda_{G}}(2\omega)\,R_{\rm in,loc}(\omega),$ (37)
and for the Gaussian regulator:
$|R_{\rm in,NL}(\omega)|\;\lesssim\;e^{-4\omega^{2}/\Lambda_{G^{2}}}\,|R_{\rm in,loc}(\omega)|.$ (38)
Even if a horizonless object would generate partial reflections in a purely local description, analytic nonlocality exponentially attenuates those reflections over the band $\omega\gtrsim\Lambda_{G}$ and smooths sharp features required for phase–coherent cavity ringing.
Combining (31) and (37) in the echo series (32), the $n$th echo carries a factor:
$\Big|R_{\rm ph,NL}(\omega)\,R_{\rm in,NL}(\omega)\Big|^{n}\;\approx\;\Big|\widetilde{K}_{\Lambda_{G}}(2\omega)\Big|^{2n}\,\Big|R_{\rm ph,loc}(\omega)\,R_{\rm in,loc}(\omega)\Big|^{n}.$ (39)
For the Gaussian form factor:
$\boxed{\Big|R_{\rm ph,NL}(\omega)\,R_{\rm in,NL}(\omega)\Big|^{n}\;\lesssim\;\exp\!\Big[-\,8n\,\omega^{2}/\Lambda_{G}^{2}\Big]\;\Big|R_{\rm ph,loc}(\omega)\,R_{\rm in,loc}(\omega)\Big|^{n}}$ (40)
which provides a direct Paley–Wiener suppression bound on the echo train. Equation (40) makes transparent that analytic nonlocality suppresses echoes both by reducing the reflectivities and by compounding the suppression with echo number $n$.
Let $\Psi_{\rm out}^{\rm loc}(\omega)$ denote the local putative echo spectrum. If the nonlocal filtering acts multiplicatively in $\omega$ with $\widetilde{K}_{\Lambda_{G}}(2\omega)$, then in the time domain the waveform is convolved with the inverse Fourier transform of the filter. For the Gaussian example:
$\Psi_{\rm out}^{\rm NL}(t)\;\simeq\;\int_{-\infty}^{+\infty}dt^{\prime}\;\frac{\Lambda_{G}}{\sqrt{16\pi}}\,\exp\!\Big[-\frac{\Lambda_{G}^{2}(t-t^{\prime})^{2}}{16}\Big]\;\Psi_{\rm out}^{\rm loc}(t^{\prime}).$ (41)
Any sharp, repeating pulses are broadened on a timescale $\Delta t\sim\Lambda_{G}^{-1}$ and reduced in peak amplitude, consistent with the frequency–space suppression (40). The same conclusion holds for general analytic entire filters. Paley–Wiener bounds imply that the filter is smooth and strongly decaying in frequency, leading to time–domain smearing and loss of echo sharpness.
The ringdown response is governed by the photon–sphere barrier and the inner boundary condition. In covariant analytic nonlocal gravity, if the remnant possesses a true horizon, the ingoing condition (34) enforces $R_{\rm in}\simeq 0$, eliminating the echo cavity. If the remnant is regular and horizonless, the appropriate inner condition is regularity (36) rather than a reflective wall, and any residual inner reflection is exponentially filtered by $\widetilde{K}_{\Lambda_{G}}(2\omega)$. In both cases, the photon–sphere reflectivity is likewise filtered, and the echo train is bounded by (40) or its general entire–function analogue.
The analytic entire–function nonlocality provides a robust mechanism to suppress ringdown echoes without requiring ad hoc dissipation or fine tuning. The suppression is a direct corollary of diffeomorphism invariance, analyticity, and Paley–Wiener bounds on the nonlocal regulator.
4 Conclusions
We have shown that gravitational–wave echo suppression is a generic and robust consequence of diffeomorphism–invariant nonlocal gravity theories constructed from analytic entire functions of the covariant d’Alembertian. The key mechanism is not model–dependent microphysics at the horizon, but rather the analyticity and Paley–Wiener bounds associated with the nonlocal regulator, which enforce exponential damping of high–frequency scattering and smear sharp structures in the effective perturbation potential.
By deriving the nonlocal master equation governing ringdown perturbations, we demonstrated that the effective photon–sphere and interior scattering amplitudes are multiplicatively filtered by the Fourier transform of the nonlocal kernel. This filtering suppresses reflection coefficients at frequencies relevant to the ringdown phase and compounds with each successive round trip between the photon sphere and the interior. As a result, any would–be echo train is exponentially attenuated with increasing frequency and echo number and is further broadened in the time domain by the intrinsic nonlocal scale.
We emphasized the distinction between regular black holes possessing true event horizons and regular horizonless compact objects. In the former case, the standard ingoing boundary condition at the horizon eliminates the inner reflection required to form a resonant cavity, independently of the ultraviolet completion. In the latter case, regularity at the center replaces the notion of a reflective surface, while analytic nonlocality strongly suppresses scattering from interior structure. In both scenarios, the two essential ingredients for observable echoes—strong inner reflectivity and a sharply defined cavity—are generically absent.
Our results have direct implications for the interpretation of current and future gravitational–wave observations. The non–detection of echoes in LIGO–Virgo–KAGRA data is fully consistent with ultraviolet–finite and quantum–gravity–motivated theories featuring regular black holes or horizonless remnants [17]. Conversely, the presence of strong, long–lived echo trains would place significant constraints on analytic nonlocal completions of gravity, favoring either non–analytic dynamics or explicit reflective boundary conditions imposed by hand.
This work highlights how fundamental properties such as diffeomorphism invariance and analyticity leave observable imprints in strong–field dynamics. Ringdown phenomenology therefore provides a complementary probe of the analytic structure and nonlocal scale of gravity, linking gravitational–wave observations to the ultraviolet completion of spacetime itself.
Acknowledgments
Research at the Perimeter Institute for Theoretical Physics is supported by the Government of Canada through Industry Canada and by the Province of Ontario through the Ministry of Research and Innovation (MRI).