### v2

# Suppression of Gravitational-Wave Echoes in Diffeomorphism-Invariant Nonlocal Quantum Gravity

Abstract
Searches for gravitational-wave echoes have been widely interpreted as probes of near-horizon structure and quantum modifications of black holes. We revisit the mechanism by which echoes are suppressed in a diffeomorphism-invariant, analytic entire-function ultraviolet completion of quantum gravity. We show that the absence of observable echoes is not due to a suppression of quasinormal-mode frequencies or a filtering of the gravitational-wave spectrum. The extreme blueshift of the proper local frequency in the near-horizon region activates the diffeomorphism-invariant entire-function regulator, which smooths out sharp reflecting inner structures and drives the reflection coefficient to zero. The nonlocal regulator acts as a covariant smearing operator on effective stress–energy and curvature, replacing sharp, partially reflective surfaces by smooth transition regions. As a result, the would-be echo cavity required for repeated reflections fails to form, and the associated phase-coherent back scattering is eliminated. This mechanism applies both to regular black holes with horizons and to horizonless compact objects generated by nonlocal smearing, and is independent of the detailed value of the nonlocal scale provided it lies at or above the Planck scale. Our results clarify the physical origin of echo suppression in nonlocal quantum gravity and demonstrate that the absence of echoes is a structural consequence of analytic nonlocality rather than a dynamical damping of gravitational-wave frequencies.
1 Introduction
Gravitational-wave observations of compact binary mergers provide a new observational window on the strong-field regime of gravity. In particular, the post-merger ringdown phase has been proposed as a sensitive probe of near-horizon or near-surface modifications of black holes, including quantum-gravity-inspired deviations from classical general relativity. A prominent phenomenological signature is the appearance of late-time gravitational-wave echoes, generated by partial reflection of perturbations between an effective potential barrier and an inner reflective structure. The absence of statistically significant echoes in current observations has therefore been used to constrain a wide class of exotic compact object models.
Ultraviolet–complete or ultraviolet–finite theories of quantum gravity can be constructed using nonlocal operators built from analytic entire functions of the covariant d’Alembertian [1, 2, 3, 4, 5, 6, 7, 8]. These theories preserve diffeomorphism invariance, avoid additional ghostlike degrees of freedom, and render quantum loop corrections finite. The nonlocality is controlled by a fundamental scale and is encoded through form factors whose analyticity implies Paley–Wiener bounds on their Fourier transforms [11, 12].
Nonlocal, analytic entire-function ultraviolet completions of quantum gravity constitute a distinct theoretical framework in which both ultraviolet finiteness and singularity resolution can be achieved without introducing new degrees of freedom or violating diffeomorphism invariance. In this approach, nonlocality enters through covariant entire functions of the d’Alembertian operator, which act as smoothing operators on stress–energy distributions and curvature invariants. In static and dynamical settings, this smearing replaces otherwise sharp geometric or material features by smooth transition regions, while preserving the classical description of spacetime at macroscopic scales.
The suppression of gravitational-wave echoes in nonlocal gravity does not originate from a selective suppression of ringdown frequencies. A crucial physical ingredient in the suppression mechanism is the invariant blueshift of the proper local frequency in the near-horizon region. For a mode of asymptotic frequency $\omega$, the locally measured frequency is given by
$\omega_{\rm loc}(r)=\frac{\omega}{\sqrt{f(r)}},$ (1)
where $f(r)=-g_{tt}(r)$. As $r\to r_{h}$ in a classical black hole geometry, $f(r)\to 0$ and $\omega_{\rm loc}\to\infty$. Even for a horizonless but highly redshifted compact object that mimics a black hole observationally, $f(r)$ becomes extremely small in the inner transition region, yielding $\omega_{\rm loc}\,\ell\gg 1$, where $\ell$ is the nonlocality length scale. It is this invariant blueshift, not a suppression of the observed gravitational-wave frequency, that activates the diffeomorphism-invariant entire-function regulator.
The suppression of ringdown echoes is a structural consequence of the smoothing of sharp reflective surfaces that would otherwise support repeated, phase-coherent backscattering. Echo formation requires the presence of nonanalytic features in the effective potential or boundary conditions, such as hard reflecting layers or discontinuous matching surfaces. The action of the entire-function regulator removes these features by replacing them with smooth, extended regions, thereby preventing the formation of a resonant echo cavity.
This mechanism operates independently of whether the underlying compact object possesses an event horizon. Both regular black holes with horizons and horizonless compact objects generated by nonlocal smearing exhibit the same qualitative behavior. The absence of sharply defined internal reflectors eliminates the conditions necessary for observable echoes. Although the nonlocal scale is taken to be Planckian, its role is not to suppress macroscopic frequencies, but to enforce analyticity and smoothness of the effective geometry. The resulting loss of phase coherence, rather than dissipative damping, is what removes the echo signal.
The suppression of gravitational–wave echoes derived in this work operates entirely at the level of classical wave propagation on a fixed background geometry. The modified Regge–Wheeler–type equations [18, 19, 20], the smearing of effective potentials, and the exponential damping of reflection coefficients follow from classical integro–differential equations of motion. However, the specific nonlocal structure responsible for these effects, namely, analytic entire functions of the covariant d’Alembertian obeying Paley–Wiener bounds is not motivated by classical considerations alone. Such analyticity conditions are imposed to ensure unitarity, ghost freedom, and ultraviolet finiteness of the quantum gravitational theory. In this sense, quantum gravity is already implicitly encoded in the classical dynamics. The same analytic properties required for a consistent quantum theory manifest classically as a frequency–space filter that suppresses sharp reflections and eliminates the resonant cavity necessary for observable echoes. The resulting ringdown phenomenology reflects quantum–gravity consistency conditions, even in the absence of explicit quantum fluctuations or loop effects.
We show that these same analytic and diffeomorphism–invariant properties have direct and robust consequences for gravitational–wave ringdown phenomenology [13, 14, 15, 16, 17].
The purpose of this paper is to provide a clear and self-consistent account of echo suppression in analytic nonlocal gravity, emphasizing its geometric origin. By doing so, we aim to place the interpretation of gravitational-wave observations in nonlocal gravity on a firmer conceptual footing and to sharpen the distinction between nonlocal smoothing effects and phenomenological damping mechanisms.
2 Regular Static Spherically Symmetric Solution
A covariant entire-function form factor $\mathcal{F}(\Box/\Lambda_{G}^{2})$ with $\mathcal{F}$ entire and $\mathcal{F}(0)=1$ can be implemented so that matter sources are nonlocally dressed, while maintaining diffeomorphism invariance [3, 4, 5]. The distributional stress–energy is replaced by an effective, smooth profile:
$T_{\mu\nu}(x)\;\longrightarrow\;T^{\rm eff}_{\mu\nu}(x):=\mathcal{F}\!\left(\frac{\Box}{\Lambda_{G}^{2}}\right)T_{\mu\nu}(x),$ (2)
where $\Box=g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}$ is the covariant d’Alembertian and $\Lambda_{G}$ is the nonlocality scale. For an exponential entire function:
$\mathcal{F}\!\left(\frac{\Box}{\Lambda_{G}^{2}}\right)=e^{\Box/\Lambda_{G}^{2}},$ (3)
the action of $\mathcal{F}$ on static sources reduces in a weakly curved, quasi-static regime to a Gaussian heat-kernel smearing in space. This is the sense in which the regulator acts as a smearing map on matter distributions.
In what follows, we derive the resulting regular black-hole and compact-object geometries in the static, spherically symmetric sector, and show explicitly the emergence of a de Sitter core and the horizon phase diagram controlled by a single dimensionless parameter.
We adopt the standard static, spherically symmetric line element:
$ds^{2}=-f(r)\,dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2},\qquad d\Omega^{2}=d\theta^{2}+\sin^{2}\theta\,d\phi^{2},$ (4)
and model the smeared source as an anisotropic but static effective fluid:
$T^{\mu}{}_{\nu}=\mathrm{diag}\!\big(-\rho(r),\,p_{r}(r),\,p_{t}(r),\,p_{t}(r)\big).$ (5)
Einstein’s equations,
$G_{\mu\nu}=8\pi G\,T_{\mu\nu},$ (6)
imply that $f(r)$ can be written in terms of the Misner–Sharp mass function $m(r)$ [21]:
$f(r)=1-\frac{2G\,m(r)}{r},\qquad m^{\prime}(r)=4\pi r^{2}\rho(r),$ (7)
where prime denotes $d/dr$. Regularity at the origin requires $m(r)\sim\mathcal{O}(r^{3})$ as $r\to 0$.
Consider a point mass $M$ localized at the origin in the distributional limit. Under the smearing map (2) with the exponential choice (3), the effective energy density is well-approximated by a Gaussian profile of width $\ell\sim\Lambda_{G}^{-1}$:
$\rho(r)=\frac{M}{(4\pi\ell^{2})^{3/2}}\exp\!\left(-\frac{r^{2}}{4\ell^{2}}\right),$ (8)
and normalized so that the total mass is $M$:
$4\pi\int_{0}^{\infty}dr\,r^{2}\rho(r)=M.$ (9)
For the horizon phase structure only the ratio $GM/\ell$ matters.
Integrating (7) with (8) gives
$m(r)$ $=4\pi\int_{0}^{r}ds\,s^{2}\rho(s)=\frac{M}{(4\pi\ell^{2})^{3/2}}\,4\pi\int_{0}^{r}ds\,s^{2}\exp\!\left(-\frac{s^{2}}{4\ell^{2}}\right)$
$=M\left[\operatorname{erf}\!\left(\frac{r}{2\ell}\right)-\frac{r}{\sqrt{\pi}\,\ell}\exp\!\left(-\frac{r^{2}}{4\ell^{2}}\right)\right],$ (10)
where $\operatorname{erf}$ is the error function. The corresponding lapse function is given by
$f(r)=1-\frac{2GM}{r}\left[\operatorname{erf}\!\left(\frac{r}{2\ell}\right)-\frac{r}{\sqrt{\pi}\,\ell}\exp\!\left(-\frac{r^{2}}{4\ell^{2}}\right)\right].$ (11)
By construction $m(r)\to M$ as $r\to\infty$, so $f(r)\to 1-2GM/r$ and the geometry is asymptotically Schwarzschild.
Near the origin, expand (10) for $r\ll\ell$:
$m(r)=\frac{4\pi}{3}\rho_{0}\,r^{3}+\mathcal{O}(r^{5}),\qquad\rho_{0}:=\rho(0)=\frac{M}{(4\pi\ell^{2})^{3/2}}.$ (12)
Then (7) yields
$f(r)=1-\frac{2G}{r}\left(\frac{4\pi}{3}\rho_{0}r^{3}\right)+\mathcal{O}(r^{4})=1-\frac{8\pi G\rho_{0}}{3}\,r^{2}+\mathcal{O}(r^{4}).$ (13)
The core is de Sitter with effective cosmological constant:
$\Lambda_{\rm eff}=8\pi G\rho_{0},\qquad f(r)=1-\frac{\Lambda_{\rm eff}}{3}r^{2}+\mathcal{O}(r^{4}),$ (14)
and all curvature invariants remain finite at $r=0$, because $\rho_{0}<\infty$ and $f(r)$ is analytic there. In particular, the would-be Schwarzschild singularity is replaced by a smooth de Sitter core. Since the Schwarzschild singularity is spacelike, its regularization necessarily replaces it with a finite spacelike de Sitter core, preserving the causal structure of the interior while eliminating curvature divergences; the resulting core is locally static but globally lacks a preferred time direction.
Horizons occur at the positive roots of $f(r)=0$. It is convenient to introduce dimensionless variables
$x:=\frac{r}{\ell},\qquad\mu:=\frac{GM}{\ell},$ (15)
so that (11) becomes
$f(x)=1-\frac{2\mu}{x}\left[\operatorname{erf}\!\left(\frac{x}{2}\right)-\frac{x}{\sqrt{\pi}}\,e^{-x^{2}/4}\right].$ (16)
The entire horizon structure is controlled by the single dimensionless parameter $\mu=GM/\ell$, equivalently, $M$ measured in units set by the nonlocal length $\ell$.
The transition between two horizons and no horizon occurs at an extremal configuration where $f$ has a double root. Denoting the extremal radius by $x_{\star}>0$, the conditions are given by
$f(x_{\star})=0,\qquad f^{\prime}(x_{\star})=0.$ (17)
From (16) these two equations implicitly determine the critical pair $(x_{\star},\mu_{\star})$:
$\mu_{\star}=\frac{x_{\star}}{2\left[\operatorname{erf}\!\left(\frac{x_{\star}}{2}\right)-\frac{x_{\star}}{\sqrt{\pi}}\,e^{-x_{\star}^{2}/4}\right]},\qquad 0=\left.\frac{d}{dx}\left\{\frac{1}{x}\left[\operatorname{erf}\!\left(\frac{x}{2}\right)-\frac{x}{\sqrt{\pi}}e^{-x^{2}/4}\right]\right\}\right|_{x=x_{\star}}.$ (18)
This provides a sharp, regulator-controlled critical point without introducing any additional dimensionless tuning.
As $\mu$ varies, one finds the standard three-regime pattern characteristic of these smeared-source solutions. $\mu>\mu_{\star}$: two distinct horizons $x_{-}$ and $x_{+}$ inner and outer. $\mu=\mu_{\star}$: one degenerate extremal horizon at $x_{\star}$. $\mu<\mu_{\star}$: no horizon a regular, horizonless compact object. Because $f(x)$ is smooth and the core is de Sitter, the $\mu<\mu_{\star}$ branch is nonsingular and geodesically extendible through $r=0$.
The effective stress tensor (5) must satisfy $\nabla_{\mu}T^{\mu}{}_{\nu}=0$. For the metric (4) this implies:
$p_{r}^{\prime}(r)+\frac{f^{\prime}(r)}{2f(r)}\big(\rho(r)+p_{r}(r)\big)+\frac{2}{r}\big(p_{r}(r)-p_{t}(r)\big)=0.$ (19)
A minimal consistent for smeared static sources is $T^{t}{}_{t}=T^{r}{}_{r}$, $p_{r}=-\rho$, which ensures regularity and simplifies (19) to a relation determining $p_{t}(r)$. The detailed microphysical interpretation of $p_{r},p_{t}$ depends on the underlying nonlocal completion, but the geometric conclusions above follow already from the single input, a smooth, conserved $T^{\rm eff}_{\mu\nu}$ generated by the entire-function smearing map.
The covariant entire-function regulator replaces distributional matter sources by smooth effective profiles, and for a smeared point mass leads to the closed-form metric (11). The geometry is asymptotically Schwarzschild, develops a de Sitter core (14), and exhibits a sharp horizon phase diagram controlled by the single parameter $\mu=GM/\ell$, with the extremal point determined implicitly by (17)–(18). This is the sense in which the regulator simultaneously enforces UV finiteness via nonlocal damping and geometrically resolves the classical Schwarzschild singularity in the static, spherically symmetric sector.
3 Suppression of Ringdown Echoes in Covariant Nonlocal Quantum Gravity
We consider the class of covariant nonlocal theories in which the gravitational action is built from curvature invariants and analytic entire functions of the covariant d’Alembertian $\Box\equiv g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}$. A representative form is given by
$S\;=\;\frac{1}{16\pi G}\int d^{4}x\,\sqrt{-g}\;\Big[R\;+\;R\,\mathcal{F}_{1}(\Box)\,R\;+\;R_{\mu\nu}\,\mathcal{F}_{2}(\Box)\,R^{\mu\nu}\;+\;C_{\mu\nu\rho\sigma}\,\mathcal{F}_{3}(\Box)\,C^{\mu\nu\rho\sigma}\Big],$ (20)
where $R$, $R_{\mu\nu}$ and $C_{\mu\nu\rho\sigma}$ denote the Ricci scalar, Ricci and Weyl curvature tensors, respectively. Each $\mathcal{F}_{i}(z)$ is analytic entire and chosen so that no additional propagating ghost poles are introduced. In ultraviolet–finite constructions one typically encounters an entire form factor or regulator of exponential type:
$\mathcal{F}(z)\;=\;\exp\!\Big[-\big(z/\Lambda_{G}^{2}\big)^{n}\Big],\qquad n\in\mathbb{Z}_{\geq 1},$ (21)
with nonlocality scale $\Lambda_{G}$. [^1] In Euclidean momentum variables $p_{E}$, one has $\mathcal{F}(p_{E}^{2})=\exp[-(p_{E}^{2}/\Lambda_{G}^{2})^{n}]$, giving strong ultraviolet damping.
We study linearized gravitational perturbations around a static, spherically symmetric regular remnant geometry describing the post–merger object:
$ds^{2}=-f(r)\,dt^{2}+\frac{dr^{2}}{g(r)}+r^{2}d\Omega^{2},\qquad f(r),g(r)\ \text{regular for}\ r\geq 0,$ (22)
allowing either a regular black hole with a true horizon at $r=r_{h}$, where $f(r_{h})=g(r_{h})=0$ or a regular horizonless compact object, where $f,g>0$ for all $r\geq 0$. We emphasize that our conclusions do not depend on a particular regular metric model; only smoothness and the analytic nonlocal structure are used.
A static black-hole spacetime does not possess spatial translation invariance and therefore admits no global momentum-space representation of the form factor. The nonlocal regulator is defined covariantly as an entire function of the d’Alembertian, $\mathcal{F}\!\left(\Box\right)$, which is well-defined on any curved background and preserves diffeomorphism invariance. In a static geometry one retains time-translation symmetry and a conserved Killing frequency. In a stationary spacetime with timelike Killing vector $\xi^{\mu}$, the conserved Killing frequency is given by
$\omega\;\equiv\;-\,k_{\mu}\,\xi^{\mu}=k_{\mu}(\partial_{t})^{\mu},$ (23)
where $k^{\mu}$ is the wave four–vector. The locally measured proper frequency by an observer with four–velocity $u^{\mu}$ is given by
$\omega_{\mathrm{loc}}\;\equiv\;-\,k_{\mu}\,u^{\mu},$ (24)
where for static observers and integral curves of $\xi^{\mu}$, the normalized four–velocity is:
$u^{\mu}\;=\;\frac{\xi^{\mu}}{\sqrt{-\,\xi^{\alpha}\xi_{\alpha}}}.$ (25)
This allows for separation of variables and reduction to a one-dimensional Regge–Wheeler scattering problem. The action of $\mathcal{F}\!\left(\Box\right)$ is understood as a quasi-local smearing over a proper length scale $\ell$, rather than as a flat-space momentum filter. Echo suppression is consequently governed by adiabatic scattering controlled by the local parameter $\omega_{loc}\ell$, and does not rely on global translation invariance.
Upon decomposing perturbations into parity sectors and spherical harmonics, the dynamics can be expressed in terms of a master variable $\Psi(t,r)$ of Regge–Wheeler-Zerilli type [18, 19, 20]. In local GR, Fourier modes $\Psi(t,r)=e^{-i\omega t}\Psi_{\omega}(r)$ satisfy:
$\Big[\partial_{r_{*}}^{2}+\omega^{2}-V_{\rm loc}(r)\Big]\Psi_{\omega}(r)=0,$ (26)
where $r_{*}$ is the tortoise coordinate defined by $dr_{*}/dr=1/\sqrt{f(r)g(r)}$, and $V_{\rm loc}(r)$ is the effective potential with a barrier near the photon sphere with $r_{ph}=3GM$.
In analytic nonlocal gravity, the operator $\Box$ acts nontrivially on perturbations and generates nonlocal modifications that can be represented as a radial convolution by a kernel whose Fourier transform is the form factor. A general and convenient parametrization of the frequency–domain equation is given by
$\Big(\partial_{r_{*}}^{2}+\omega^{2}\Big)\Psi_{\omega}(r_{*})\;=\;\int_{-\infty}^{+\infty}dr_{*}^{\prime}\;K_{\Lambda_{G}}(r_{*}-r_{*}^{\prime})\,V_{\rm loc}(r_{*}^{\prime})\,\Psi_{\omega}(r_{*}^{\prime})\,$, (27)
where $K_{\Lambda_{G}}$ is a smooth kernel of width $\Delta r_{*}\sim\Lambda_{G}^{-1}$ in the local proper measure, determined by $\mathcal{F}$. Equation (27) states that the nonlocality smears the potential and modifies scattering. We define the smeared effective potential:
$V_{\rm eff}(r_{*})\;\equiv\;\int_{-\infty}^{+\infty}dr_{*}^{\prime}\;K_{\Lambda_{G}}(r_{*}-r_{*}^{\prime})\,V_{\rm loc}(r_{*}^{\prime})\,$, (28)
To make this explicit, consider scattering at frequency $\omega$ from a localized potential barrier. In the Born approximation, the complex reflection amplitude in the local problem (26) is given by
$R_{\rm loc}(\omega)\;\simeq\;\frac{1}{2i\omega}\int_{-\infty}^{+\infty}dr_{*}\;e^{2i\omega r_{*}}\,V_{\rm loc}(r_{*}).$ (29)
Replacing $V_{\rm loc}$ by the smeared $V_{\rm eff}=K_{\Lambda_{G}}*V_{\rm loc}$, the nonlocal reflection amplitude becomes:
$R_{\rm NL}(\omega)$ $\simeq\frac{1}{2i\omega}\int dr_{*}\;e^{2i\omega r_{*}}\,V_{\rm eff}(r_{*})$
$=\frac{1}{2i\omega}\int dr_{*}^{\prime}\;V_{\rm loc}(r_{*}^{\prime})e^{2i\omega r_{*}^{\prime}}\int dx\;K_{\Lambda_{G}(x)}e^{2i\omega x}$
$=\widetilde{K}_{\Lambda_{G}}(2\omega)\,R_{\rm loc}(\omega).$ (30)
The analytic nonlocality multiplicatively suppresses reflection through the factor $\widetilde{K}_{\Lambda_{G}}(2\omega)$. For the Gaussian filter:
$|R_{\rm NL}(\omega)|\;\simeq\;e^{-4\omega^{2}/\Lambda_{G}^{2}}\,|R_{\rm loc}(\omega)|.$ (31)
This is the basic Paley–Wiener suppression mechanism [11, 12]. The same analyticity that yields ultraviolet finiteness forces reflection coefficients to be exponentially damped at frequencies above the nonlocal scale.
The photon–sphere barrier reflectivity $R_{\rm ph}(\omega)$ entering ringdown scattering is therefore filtered as:
$R_{\rm ph,NL}(\omega)\;\approx\;\widetilde{K}_{\Lambda_{G}}(2\omega)\,R_{\rm ph,loc}(\omega).$ (32)
Echo trains are commonly modeled as repeated partial reflections between the photon–sphere barrier and an inner reflector. In frequency space one may write the outgoing waveform as:
$\Psi_{\rm out}(\omega)\;\propto\;T_{\rm ph}(\omega)\,\Psi_{\rm in}(\omega)\,\sum_{n=0}^{\infty}\Big[R_{\rm ph}(\omega)\,R_{\rm in}(\omega)\,e^{2i\omega\Delta t}\Big]^{n},$ (33)
where $T_{\mathrm{ph}}(\omega)$ denotes the complex transmission coefficient for a gravitational perturbation mode of frequency $\omega$ propagating through the photon-sphere and angular-momentum barrier in the Regge–Wheeler potential. $R_{\rm in}(\omega)$ is the inner reflection amplitude and $\Delta t$ is the round–trip time between the photon sphere and the inner scattering region expressed in tortoise coordinate. The transfer function is given by
$\mathcal{G}(\omega)\;\equiv\;\frac{1}{1-R_{\rm ph}(\omega)\,R_{\rm in}(\omega)\,e^{2i\omega\Delta t}}.$ (34)
The existence of observable echoes requires two conditions, a non-negligible inner reflectivity $|R_{\rm in}(\omega)|\not\ll 1$ over the relevant band, and sufficiently sharp scattering phase coherence to sustain repeated returns, so that $|R_{\rm ph}R_{\rm in}|$ is not rapidly suppressed with $\omega$. We now show that analytic entire–function nonlocality generically violates both conditions.
If the remnant has a true horizon at $r=r_{h}$, then $r_{*}\to$-$\infty$ as $r\to r_{h}^{+}$ and the physical boundary condition for perturbations is purely ingoing at the horizon:
$\Psi_{\omega}(r_{*})\;\sim\;e^{-i\omega r_{*}},\qquad r_{*}\to-\infty.$ (35)
This implies the absence of an outgoing component from the interior:
$R_{\rm in}(\omega)\;\simeq\;0\qquad\Rightarrow\qquad\mathcal{G}(\omega)\approx 1.$ (36)
A regular black hole with a horizon does not support an echo cavity, irrespective of the ultraviolet completion, provided the completion does not introduce a reflective surface outside the horizon. Analytic entire–function nonlocality, being smooth and diffeomorphism invariant, does not generate such a hard surface; rather it smears curvature and sources.
For a regular horizonless compact object, $r_{*}$ remains finite throughout the interior and the appropriate inner condition is regularity at the center. For typical parity sectors one has near $r=0$:
$\Psi_{\omega}(r)\;\propto\;r^{\ell+1}\quad(r\to 0),$ (37)
which is not equivalent to imposing a reflective wall. Any effective inner reflection arises instead from smooth interior structure turning points in $V_{\rm eff}$, and is therefore intrinsically sensitive to the nonlocal smearing.
Applying the same reasoning as in (30), the inner reflectivity is filtered:
$R_{\rm in,NL}(\omega)\;\approx\;\widetilde{K}_{\Lambda_{G}}(2\omega)\,R_{\rm in,loc}(\omega),$ (38)
and for the Gaussian regulator:
$|R_{\rm in,NL}(\omega)|\;\lesssim\;e^{-4\omega^{2}/\Lambda_{G^{2}}}\,|R_{\rm in,loc}(\omega)|.$ (39)
Even if a horizonless object would generate partial reflections in a purely local description, analytic nonlocality exponentially attenuates those reflections over the band $\omega\gtrsim\Lambda_{G}$ and smooths sharp features required for phase–coherent cavity ringing.
Combining (38) in the echo series (33), the $n$th echo carries a factor:
$\Big|R_{\rm ph,NL}(\omega)\,R_{\rm in,NL}(\omega)\Big|^{n}\;\approx\;\Big|\widetilde{K}_{\Lambda_{G}}(2\omega)\Big|^{2n}\,\Big|R_{\rm ph,loc}(\omega)\,R_{\rm in,loc}(\omega)\Big|^{n}.$ (40)
For the Gaussian form factor:
$\Big|R_{\rm ph,NL}(\omega)\,R_{\rm in,NL}(\omega)\Big|^{n}\;\lesssim\;\exp\!\Big[-\,8n\,\omega^{2}/\Lambda_{G}^{2}\Big]\;\Big|R_{\rm ph,loc}(\omega)\,R_{\rm in,loc}(\omega)\Big|^{n},$ (41)
which provides a direct Paley–Wiener suppression bound on the echo train. Equation (41) makes transparent that analytic nonlocality suppresses echoes both by reducing the reflectivities and by compounding the suppression with echo number $n$.
Let $\Psi_{\rm out}^{\rm loc}(\omega)$ denote the local putative echo spectrum. If the nonlocal filtering acts multiplicatively in $\omega$ with $\widetilde{K}_{\Lambda_{G}}(2\omega)$, then in the time domain the waveform is convolved with the inverse Fourier transform of the filter. For the Gaussian example:
$\Psi_{\rm out}^{\rm NL}(t)\;\simeq\;\int_{-\infty}^{+\infty}dt^{\prime}\;\frac{\Lambda_{G}}{\sqrt{16\pi}}\,\exp\!\Big[-\frac{\Lambda_{G}^{2}(t-t^{\prime})^{2}}{16}\Big]\;\Psi_{\rm out}^{\rm loc}(t^{\prime}).$ (42)
Any sharp, repeating pulses are broadened on a timescale $\Delta t\sim\Lambda_{G}^{-1}$ and reduced in peak amplitude, consistent with the frequency–space suppression (41). The same conclusion holds for general analytic entire filters. Paley–Wiener bounds imply that the filter is smooth and strongly decaying in frequency, leading to time–domain smearing and loss of echo sharpness.
The ringdown response is governed by the photon–sphere barrier and the inner boundary condition. In covariant analytic nonlocal gravity, if the remnant possesses a true horizon, the ingoing condition (35) enforces $R_{\rm in}\simeq 0$, eliminating the echo cavity. If the remnant is regular and horizonless, the appropriate inner condition is regularity (37) rather than a reflective wall, and any residual inner reflection is exponentially filtered by $\widetilde{K}_{\Lambda_{G}}(2\omega)$. In both cases, the photon–sphere reflectivity is likewise filtered, and the echo train is bounded by (41) or its general entire–function analogue.
Let us consider the role of the Paley–Wiener entire-function factor:
$\mathcal{F}(\omega)\equiv\exp\!\left(-\frac{\omega^{2}}{\Lambda_{G}^{2}}\right),$ (43)
which appears in nonlocal response functions and in the frequency-domain matching of Regge–Wheeler (RW) modes. For astrophysical ringdown, the relevant frequencies are set by the black-hole scale:
$\omega_{\rm QNM}\sim\mathcal{O}\!\left(\frac{1}{M}\right),\qquad f_{\rm QNM}\sim\mathcal{O}\!\left(\frac{1}{2\pi M}\right).$ (44)
For $M\sim 10\,M_{\odot}$ this corresponds to $f_{\rm QNM}\sim 10^{2}$$10^{3}\,$Hz. If $\Lambda_{G}$ is Planckian or even many orders below Planck, then $\omega_{\rm QNM}\ll\Lambda_{G}$ and we obtain:
$\mathcal{F}(\omega_{\rm QNM})=\exp\!\left(-\frac{\omega_{\rm QNM}^{2}}{\Lambda_{G}^{2}}\right)=1-\mathcal{O}\!\left(\frac{\omega_{\rm QNM}^{2}}{\Lambda_{G}^{2}}\right)\simeq 1.$ (45)
This explains why one should not expect any direct suppression of the black-hole frequency itself: the nonlocal scale is so high compared to $\omega_{\rm QNM}$ that (43) is essentially unity throughout the LIGO/Virgo/KAGRA band.
Echo phenomenology depends not on suppressing $\omega_{\rm QNM}$, but on the existence of a hard, highly reflective structure near the would-be horizon. In the RW problem, echoes arise only if there is an additional inner reflection coefficient $\mathcal{R}_{\rm in}(\omega)$ of magnitude close to unity, creating a cavity between the photon-sphere barrier and the inner surface. The observable echo train is controlled schematically by
$\tilde{Z}_{\rm echo}(\omega)\;\propto\;\frac{\mathcal{T}_{\rm ps}(\omega)\,\mathcal{R}_{\rm in}(\omega)\,e^{2i\omega L}}{1-\mathcal{R}_{\rm ps}(\omega)\,\mathcal{R}_{\rm in}(\omega)\,e^{2i\omega L}},$ (46)
where $L$ is the cavity length in tortoise coordinate, and $\mathcal{R}_{\rm ps},\mathcal{T}_{\rm ps}$ are the photon-sphere reflection and transmission amplitudes. The necessary condition for echoes is given by
$|\mathcal{R}_{\rm in}(\omega)|\not\ll 1\quad\text{over the ringdown band}.$ (47)
The nonlocal effect is that entire-function UV completion generically replaces a sharp distributional interaction localized at a surface by a smeared interaction of width $\ell\equiv\Lambda_{G}^{-1}$, so that would-be reflecting walls or abrupt transitions in effective medium properties are smoothed over a proper length $\sim\ell$. A convenient RW model is to represent the inner structure by an effective potential contribution localized near $r_{*}=r_{*0}$:
$V_{\rm in}(r_{*})=\lambda\,\delta(r_{*}-r_{*0})\qquad\longrightarrow\qquad V_{\rm in}^{\rm NL}(r_{*})=\lambda\,\frac{1}{\sqrt{\pi}\,\ell}\,\exp\!\left[-\frac{(r_{*}-r_{*0})^{2}}{\ell^{2}}\right],$ (48)
which is the action of an entire-function smearing map on a distributional source. For a monochromatic RW mode $\psi_{\omega}(r_{*})\sim e^{-i\omega t}$, the reflection amplitude off a localized perturbation is proportional to the Fourier transform of the profile. Consequently, the smearing in (48) yields the universal frequency dependence:
$\mathcal{R}_{\rm in}^{\rm NL}(\omega)\simeq\mathcal{R}_{\rm in}^{(0)}(\omega)\,\exp\!\left(-\frac{\omega^{2}\ell^{2}}{4}\right)=\mathcal{R}_{\rm in}^{(0)}(\omega)\,\exp\!\left(-\frac{\omega^{2}}{4\Lambda_{G}^{2}}\right),$ (49)
up to model-dependent phases and prefactors, where $\mathcal{R}_{\rm in}^{(0)}$ is the unsmeared local inner reflectivity. Equation (49) shows the correct physical separation of roles. The nonlocal factor does not suppress the ringdown frequency $\omega_{\rm QNM}$ itself as demonstrate in (45). Instead, it suppresses the ability of a sharp surface to reflect RW modes, by exponentially damping of the proper local frequency $\omega_{loc}$ components associated with an abrupt impedance mismatch. In other words, nonlocality eliminates the hard reflecting boundary condition by turning it into a smooth kernel.
4 Regge–Wheeler Matching and Vanishing of the Inner Reflection Coefficient
The frequency domain, $\Psi_{\omega\ell}(r_{*})$ is governed by the Regge–Wheeler equation:
$\left[\frac{d^{2}}{dr_{*}^{2}}+\omega^{2}-V_{\rm eff}(r)\right]\Psi_{\omega\ell}(r_{*})=0,\qquad\frac{dr_{*}}{dr}=f^{-1}(r),$ (50)
where $f(r)$ is the background lapse function and $V_{\rm eff}(r)$ is the standard axial or polar effective potential. In models that admit gravitational-wave echoes, an effective inner boundary condition is imposed at some $r_{*}=r_{*}^{(0)}$:
$\Psi_{\omega\ell}(r_{*})\;\xrightarrow[r_{*}\to r_{*}^{(0)}]{}e^{-i\omega r_{*}}+R_{\rm in}(\omega)\,e^{+i\omega r_{*}},$ (51)
where $R_{\rm in}(\omega)$ represents the inner reflection coefficient. A nonvanishing $R_{\rm in}$ signals the presence of a sharp reflective structure capable of generating late-time echoes.
In analytic nonlocal gravity, the dynamics is modified by an entire-function form factor involving the covariant d’Alembertian. At the level of the one-dimensional scattering problem, this implies that sharp structures in the effective potential are replaced by a smeared, analytic profile,
$V_{\rm eff}(r_{*})\;\longrightarrow\;V_{\rm eff}^{\rm NL}(r_{*})\equiv e^{\Box/\Lambda_{G}^{2}}\,V_{\rm eff}(r_{*}).$ (52)
In a locally flat approximation across the narrow inner transition region, the action of the nonlocal operator may be written as a convolution:
$V_{\rm eff}^{\rm NL}(r_{*})=\int_{-\infty}^{+\infty}ds\;\frac{1}{\sqrt{4\pi}\,\ell}\,e^{-s^{2}/(4\ell^{2})}\,V_{\rm eff}(r_{*}-s),\qquad\ell\equiv\Lambda_{G}^{-1}.$ (53)
Equation (53) shows explicitly that any nonanalytic feature—such as a thin shell, hard surface, or discontinuous matching layer—is replaced by a smooth transition region of width $\mathcal{O}(\ell)$.
To extract the inner reflection coefficient generated by the smooth transition, we rewrite Eq. (50) in Lippmann–Schwinger form:
$\Psi_{\omega\ell}(r_{*})=e^{-i\omega r_{*}}+\int_{-\infty}^{+\infty}dr_{*}^{\prime}\;G_{\omega}(r_{*}-r_{*}^{\prime})\,V_{\rm eff}^{\rm NL}(r_{*}^{\prime})\,\Psi_{\omega\ell}(r_{*}^{\prime}),$ (54)
where the one-dimensional Green function is given by
$G_{\omega}(x)=\frac{i}{2\omega}\,e^{i\omega|x|}.$ (55)
In the Born approximation, sufficient to determine the scaling of the reflected amplitude, one substitutes $\Psi_{\omega\ell}\to e^{-i\omega r_{*}}$ in the integral. The coefficient of the outgoing wave $e^{+i\omega r_{*}}$ on the inner side then yields:
$R_{\rm in}(\omega)=\frac{1}{2i\omega}\int_{-\infty}^{+\infty}dr_{*}\;e^{2i\omega r_{*}}\,V_{\rm eff}^{\rm NL}(r_{*}).$ (56)
Thus, $R_{\rm in}(\omega)$ is proportional to the Fourier transform of the inner transition profile evaluated at wavenumber $2\omega$. Reflection therefore directly probes the sharpness of the effective potential.
To make this explicit, consider a localized inner structure whose nonlocal image is well approximated by an analytic profile:
$V_{\rm eff}^{\rm NL}(r_{*})\simeq V_{0}\,\exp\!\left[-\frac{(r_{*}-r_{*}^{(0)})^{2}}{\ell^{2}}\right].$ (57)
Substitution into Eq. (56) yields:
$R_{\rm in}(\omega)=\frac{V_{0}}{2i\omega}\,\sqrt{\pi}\,\ell\,e^{2i\omega r_{*}^{(0)}}\,\exp\!\left[-\omega^{2}\ell^{2}\right].$ (58)
The exponential factor in Eq. (58) arises solely from the analyticity of the smeared potential and represents suppression of the reflection coefficient, not suppression of the physical ringdown frequencies.
Although $\ell$ is Planckian, the relevant quantity controlling reflection is the local proper frequency. Near a would-be horizon:
$\omega_{\rm loc}(r)=\frac{\omega}{\sqrt{f(r)}}\;\xrightarrow[r\to r_{h}$$\infty$. (59)
Expressed in terms of proper distance across the smooth transition region, Eq. (58) implies
$|R_{\rm in}(\omega)|\;\sim\;\exp\!\left[-\omega_{\rm loc}^{2}\,\ell^{2}\right]\;\longrightarrow\;0.$ (60)
Therefore, the inner reflection coefficient vanishes because the entire-function regulator removes sharp reflective structures from the effective geometry. The absence of echoes follows from the loss of phase-coherent inner reflection, not from any suppression or filtering of black-hole quasinormal-mode frequencies.
The quantity that controls reflection from the inner transition region is not the asymptotic frequency $\omega$ measured by a distant observer, but the local proper frequency measured by a physical observer at radius $r$. For a static, spherically symmetric background:
$ds^{2}=-f(r)\,dt^{2}+f^{-1}(r)\,dr^{2}+r^{2}d\Omega^{2},$ (61)
a perturbative mode with time dependence $\Psi\sim e^{-i\omega t}$ has locally measured frequency:
$\omega_{\rm loc}(r)=\frac{\omega}{\sqrt{f(r)}}.$ (62)
This is the standard gravitational blueshift of frequencies toward regions of large redshift.
The relevance of $\omega_{\rm loc}$ follows from the fact that mode conversion between ingoing and outgoing solutions is governed by the ratio of the local oscillation scale to the length scale over which the effective potential varies. In analytic nonlocal gravity, the entire-function regulator smooths any would-be inner reflective structure over a proper length $\ell\equiv\Lambda_{G}^{-1}$, so that the dimensionless parameter controlling reflection is $\omega_{\rm loc}\ell$, rather than $\omega\ell$.
For a black hole with an event horizon at $r=r_{h}$, one has $f(r_{h})=0$, and it follows:
$\omega_{\rm loc}(r)=\frac{\omega}{\sqrt{f(r)}}\;\xrightarrow[r\to r_{h}$$\infty$. (63)
As a result, even gravitational waves with astrophysically small asymptotic frequencies acquire arbitrarily large local proper frequencies in the vicinity of the horizon. Since the inner transition region is smoothed on a scale $\ell$, this implies:
$\omega_{\rm loc}\,\ell\gg 1,$ (64)
and therefore an exponential suppression of the inner reflection coefficient.
The same mechanism operates in regular, horizonless compact objects generated by nonlocal smearing. Although $f(r)$ does not vanish, such geometries typically contain a deep redshift well, characterized by a minimum value $f_{\rm min}\ll 1$ in the inner transition region. The corresponding maximal local frequency is given by
$\omega_{\rm loc}^{\rm max}=\frac{\omega}{\sqrt{f_{\rm min}}},$ (65)
which can be enormously larger than $\omega$ even for moderate redshifts. Consequently, for Planckian smoothing lengths $\ell\sim\ell_{\rm Pl}$ one again finds:
$\omega_{\rm loc}^{\rm max}\,\ell\gg 1,$ (66)
ensuring that the inner transition region is traversed adiabatically and that ingoing and outgoing modes do not mix.
In both horizon and horizonless cases, the limit $\omega_{\rm loc}\ell\gg 1$ implies that the effective geometry is impedance matched across the inner transition region, yielding:
$R_{\rm in}(\omega)\;\to\;0.$ (67)
The suppression of gravitational-wave echoes therefore arises from the loss of phase-coherent inner reflection enforced by nonlocal smoothing, rather than from any suppression or filtering of the physical ringdown frequencies themselves.
This is why the same UV completion can leave classical Kerr-Schwarzschild ringdown essentially unchanged, while still predicting no observable echoes from regular and horizonless compact objects. Echoes require $|\mathcal{R}_{\rm in}|\sim 1$, whereas entire-function smearing drives $|\mathcal{R}_{\rm in}^{\rm NL}(\omega)|\ll 1$ across the relevant band whenever the would-be surface is not fundamentally sharp. Operationally, the nonlocal theory replaces a Dirichlet-Neumann-type inner boundary condition at $r_{*0}$ by a nonlocal relation of the form:
$\psi_{\omega}(r_{*0})\;\;\longrightarrow\;\;\int_{-\infty}^{+\infty}\!{\rm d}r^{\prime}_{*}\;K_{\ell}(r_{*0}-r^{\prime}_{*})\,\psi_{\omega}(r^{\prime}_{*}),\qquad K_{\ell}(x)=\frac{1}{\sqrt{\pi}\,\ell}\,e^{-x^{2}/\ell^{2}},$ (68)
which removes the hard wall responsible for large reflection in the local horizonless models.
The cavity phase $e^{2i\omega L}$ in (46) is unaffected at leading order when $\omega\ll\Lambda_{G}$; the suppression mechanism acts through $\mathcal{R}_{\rm in}(\omega)$, not by shifting the characteristic black-hole frequencies. The nonlocality does not suppress the black-hole ringdown frequency, but it suppresses the reflectivity of any near-horizon or horizonless surface by smoothing it over the nonlocal length $\ell$, thereby eliminating observable echoes.
The suppression of echoes in the present framework follows from two independent but physically linked ingredients: (i) smoothing of the inner transition region by the nonlocal entire-function regulator over a proper length scale $\ell$, and (ii) the extreme gravitational redshift required for black-hole mimicry.
Consider a horizonless compact object whose surface or inner transition region lies at:
$R=2M(1+\epsilon),\qquad 0<\epsilon\ll 1.$ (69)
The exterior metric function satisfies:
$f(R)=1-\frac{2M}{R}=\frac{\epsilon}{1+\epsilon}\simeq\epsilon,$ (70)
so that the gravitational redshift at the surface is given by
$1+z=\frac{1}{\sqrt{f(R)}}\simeq\epsilon^{-1/2}.$ (71)
An object that is observationally indistinguishable from a classical black hole must have $\epsilon\ll 1$ in order to suppress detectable surface emission and luminosity from infalling matter. Extreme redshift is therefore not an additional assumption but a necessary condition for black-hole mimicry.
A perturbation mode of Killing frequency $\omega$ measured at infinity is locally measured in the inner transition region with proper frequency:
$\omega_{\rm loc}(R)=\frac{\omega}{\sqrt{f(R)}}\simeq\frac{\omega}{\sqrt{\epsilon}}.$ (72)
Even for modest astrophysical ringdown frequencies $\omega$, the locally measured frequency becomes parametrically large as $\epsilon\to 0$.
The reflection coefficient from a smooth transition region of proper thickness $\ell$ depends on the dimensionless adiabaticity parameter:
$x\equiv\omega_{\rm loc}\,\ell=\frac{\omega\ell}{\sqrt{\epsilon}}.$ (73)
When $x\gg 1$, the mode probes the transition region in the adiabatic regime and the inner reflection coefficient is strongly suppressed:
$R_{\rm in}(\omega)\to 0.$ (74)
Ultra-compact horizonless configurations that genuinely mimic black holes have $\epsilon\ll 1$. The large gravitational blueshift ensures $\omega_{\rm loc}\ell\gg 1$ even if $\omega\ell\ll 1$ at infinity. Echo suppression therefore arises from the combination of extreme redshift and nonlocal smoothing of sharp reflective structures, rather than from any suppression of the asymptotic ringdown frequency itself.
The analytic entire–function nonlocality provides a robust mechanism to suppress ringdown echoes without requiring ad hoc dissipation or fine tuning. The suppression is a direct corollary of diffeomorphism invariance, analyticity, and Paley–Wiener bounds on the nonlocal regulator.
Conclusions
We have demonstrated that diffeomorphism-invariant nonlocal gravity regulated by an analytic entire-function form factor does not produce observable gravitational-wave echoes in the merger ringdown phase of compact objects that are asymptotically Schwarzschild or Kerr. The suppression of echoes does not arise from any modification of the asymptotic quasinormal-mode spectrum, nor from a filtering of the observed gravitational-wave frequency. The ringdown frequencies measured at infinity remain those determined by the classical exterior geometry.
The essential physical mechanism is geometric. For a mode of asymptotic frequency $\omega$, the invariant proper local frequency is $\omega_{\rm loc}(r)=\frac{\omega}{\sqrt{f(r)}}$, where $f(r)=-g_{tt}(r)$. As $r\to r_{h}$ in a classical black-hole geometry, $f(r)\to 0$ and $\omega_{\rm loc}\to\infty$. Even for a horizonless but strongly redshifted compact object that mimics a black hole observationally, $f(r)$ becomes extremely small in the inner transition region, yielding $\omega_{\rm loc}\,\ell\gg 1$, where $\ell$ is the nonlocality length scale. This invariant blueshift activates the entire-function regulator in the near-horizon region. The regulator smooths out sharp inner transition layers that would otherwise act as partially reflecting boundaries. As a consequence, the inner reflection coefficient satisfies $R_{\rm in}\to 0$, eliminating the physical mechanism required for echo generation.
Echo suppression is therefore a consequence of nonlocal geometric smoothing induced by large invariant local frequencies, rather than any ad hoc frequency cutoff or damping of the observed waveform. Although the nonlocality scale is of order the Planck length, the near-horizon blueshift renders the invariant argument of the entire-function form factor large, even for low-frequency modes measured at infinity.
Within this framework, compact objects consistent with the classical exterior solutions of general relativity do not produce detectable ringdown echoes, even when the interior is rendered regular by nonlocal ultraviolet completion.
Our results highlight a conceptual distinction between local exotic compact object models and nonlocal quantum gravity. In local theories, echoes are a generic consequence of imposing hard boundary conditions near the would-be horizon. In contrast, in an analytic entire-function completion of gravity, such boundary conditions are not physically admissible. They are automatically smoothed by the same nonlocality that renders the theory ultraviolet finite and free of curvature singularities. The absence of observable echoes emerges as a prediction of nonlocal quantum gravity, rather than a fine-tuned or phenomenological assumption.
Future high-precision ringdown observations will continue to test this picture by constraining deviations from the classical quasinormal spectrum. Within the framework studied here, any confirmed detection of sharp, high-amplitude echoes would point not to a different choice of nonlocal scale, but to the breakdown of the entire-function nonlocal completion itself. In this sense, gravitational-wave echoes provide a clean observational handle on the fundamental admissibility of sharp structures in quantum gravity.
Acknowledgments
I thank Ethan Thompson for helpful discussions. Research at the Perimeter Institute for Theoretical Physics is supported by the Government of Canada through Industry Canada and by the Province of Ontario through the Ministry of Research and Innovation (MRI).